# A higher-dimensional version of Fáry’s Theorem

### v2

@@ Abstract.
We prove a generalization of István Fáry’s celebrated theorem to higher dimensions. Namely, we show that if a finite simplicial complex $X$ can be a piece-wise linearly embedded into a $d$-dimensional PL manifold $M$, then there is a triangulation of $M$ containing $X$ as a subcomplex.
April 26, 2024
@@
Choose now a ”bad” homeomorphism $\varphi$ sending $(\sigma,\partial\sigma)$ to the pair $(n\ast\gamma,\gamma)$ so that in particular some interior point of $\sigma$ is mapped to $\gamma$. Obviously, every point of the image $\varphi(\sigma^{\circ})$ apart from $n$ has a flat neighborhood with respect to $\sigma$, that is, a neighborhood where the embedding is homeomorphic to the standard embedding of $\mathbb{R}^{2}$ into $\mathbb{R}^{4}$.
@@
Lemma 4 (Section I.2, [Bing83]).
@@ 1. The proof
Consider now the biased derived subdivision of a pair of simplicial complexes $\Gamma\subseteq\Delta$: It is obtained by applying stellar subdivisions at faces of $\Delta$ not in $\Gamma$, in order of reverse inclusion; for illustration see Figure 1.2 (left). In the followig lemma we relate it to strongly induced subcomplexes: A subcomplex $\Gamma$ of a simplicial complex $\Delta$ is a strongly induced subcomplex of $\Delta$ if for every simplex $\sigma$ in $\Delta\setminus\Gamma$, $\Gamma\cap\mathrm{st}_{\sigma}\,\Delta$ consists of a single simplex, where $\mathrm{st}_{\sigma}\,\Delta$ is the star of $\sigma$ in $\Delta$, that is, the simplicial closure of all faces of $\Delta$ containing $\sigma$. Note that strongly induced subcomplex is automatically an induced subcomplex, the opposite implication does not hold as demonstrated on Figure 1.1.
@@ 1.2. Edge contractions & subdivisions
Recall that by results of Alexander [Ale30], and Newman [New31, Corrolary 10:2d], two PL homeomorphic complexes are related by edge subdivisions and valid edge contractions.
@@
Note that Lemmata 7 and 8 guarantee that $M^{\prime}$ is indeed a triangulation, meaning that it stays a simplicial complex in each step during the proccess (in order to extend the edge contraction from a subcomplex to the whole complex we need the subcomplex to be strongly induced, see Lemma 7). Note also that the biased derived subdivisions following edge contractions used in Lemma 8 do not change the respective subcomplex, so they are irrelevant with respect to the sequence of edge subdivisions and valid edge contractions we need to perform (by the result of Alexander and Newman), however they are obviously relevant for the triangulation. ∎
Acknowledgements
@@
The first author is supported by the ISF, the Centre National de Recherche Scientifique, and the Horizon Europe ERC Grant number: 101045750 / Project acronym: HodgeGeoComb. The second author is supported by the GAČR grant no. 22-19073S.

### v3

@@ Abstract.
We prove a generalization of István Fáry’s celebrated theorem to higher dimensions. Namely, we show that if a finite simplicial complex $X$ can be piecewise linearly embedded into a $d$-dimensional PL manifold $M$, then there is a triangulation of $M$ containing $X$ as a subcomplex.
March 11, 2025
@@
Choose now a ”bad” homeomorphism $\varphi$ sending $(\sigma,\partial\sigma)$ to the pair $(n\ast\gamma,\gamma)$ so that in particular some interior point of $\sigma$ is mapped to $n$. Obviously, every point of the image $\varphi(\sigma^{\circ})$ apart from $n$ has a flat neighborhood with respect to $\sigma$, that is, a neighborhood where the embedding is homeomorphic to the standard embedding of $\mathbb{R}^{2}$ into $\mathbb{R}^{4}$.
@@
Lemma 4 (Section I 2[Bing83]).
@@ 1. The proof
Consider now the biased derived subdivision of a pair of simplicial complexes $\Gamma\subseteq\Delta$: It is obtained by applying stellar subdivisions at faces of $\Delta$ not in $\Gamma$, in order of reverse inclusion; for illustration see Figure 1.2 (left). In the following lemma we relate it to strongly induced subcomplexes: A subcomplex $\Gamma$ of a simplicial complex $\Delta$ is a strongly induced subcomplex of $\Delta$ if for every simplex $\sigma$ in $\Delta\setminus\Gamma$, $\Gamma\cap\mathrm{st}_{\sigma}\,\Delta$ consists of a single simplex, where $\mathrm{st}_{\sigma}\,\Delta$ is the star of $\sigma$ in $\Delta$, that is, the simplicial closure of all faces of $\Delta$ containing $\sigma$. Note that strongly induced subcomplex is automatically an induced subcomplex, the opposite implication does not hold as demonstrated on Figure 1.1.
@@ 1.2. Edge contractions & subdivisions
Recall that by results of Alexander [Ale30, Corrolary 10:2d], and Newman [New31], two PL homeomorphic complexes are related by edge subdivisions and valid edge contractions.
@@
Note that Lemmata 7 and 8 guarantee that $M^{\prime}$ is indeed a triangulation, meaning that it stays a simplicial complex in each step during the process (in order to extend the edge contraction from a subcomplex to the whole complex we need the subcomplex to be strongly induced, see Lemma 7). Note also that the biased derived subdivisions following edge contractions used in Lemma 8 do not change the respective subcomplex, so they are irrelevant with respect to the sequence of edge subdivisions and valid edge contractions we need to perform (by the result of Alexander and Newman), however they are obviously relevant for the triangulation. ∎
Acknowledgments
@@
The first author is supported by the Centre National de Recherche Scientifique, and the Horizon Europe ERC Grant number: 101045750 / Project acronym: HodgeGeoComb. The second author is supported by the GAČR grant no. 22-19073S.